The slope of a linear equation is 4/5 the intercept (0,6) what is the equation
Answer by tarungehlot(9)

m= 4/5(given)
let
y= mx +c
be equation of line
y= 4/5 x + 6
if put now x=0
we gets
intercept y= 6=c
or (0,6)
hence we are right